At 200.0 K, a gas fills a 475.0 mL balloon. What will be the volume of the balloon if the temperature is raised by 60.0 K?


1
Expert's answer
2022-02-02T16:46:05-0500

T1 = 200.0 K

V1 = 475.0 mL

P1 = P2 = const

T2 = 200.0 K + 60.0 K = 260.0 K

V2 = unknown


Solution:

Since the pressure doesn't change, Charles' Law can be used.

Charles' Law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the new volume, solve the equation for V2:

V2 = V1T2 / T1

V2 = (475.0 mL × 260.0 K) / (200.0 K) = 617.5 mL

V2 = 617.5 mL


Answer: The volume of the balloon will be 617.5 mL